OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = x + (x - 6) What is the regular price of a child's ticket? My answer is $56 Am I right?

OpenStudy (anonymous):

\[78 = \frac{ x }{ 2 } + 2(\frac{ x - 6 }{ 2 }) = \frac{ x }{ 2 } + (x - 6)\] So, x = 56 Your answer is right, but your equation was not written correctly.

OpenStudy (anonymous):

In short, that first x was supposed to be written x/2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

you're welcome!